# A Hamiltonian for 1:1 Rotational Secondary Resonances, and Application to Small Satellites of Saturn and Jupiter

### v1

@@ Abstract
In this work, we study the dynamics of rotation of the small satellites Methone and Aegaeon and revisit previous works on the rotation of Prometheus, Metis, and Amalthea. In all cases, the surfaces of section computed with the standard spin-orbit model reveal that the synchronous regime shares another large domain in the rotation phase space. We reproduce and apply the hamiltonian theory given in Wisdom (2004) to analytically characterize the detected structure as being a secondary resonance where the period of oscillations around the synchronism is similar to the orbital period. Being that the current rotational states of this sort of satellite should be synchronous (Thomas and Helfenstein 2020), our results can be taken into account in evolutionary studies of their rotation.
@@ 1 Introduction
where $\theta$ is an angle of rotation of the satellite measured from an inertial line; $p=C\dot{\theta}$ is the angular momentum of rotation around the $z-$axis assumed to be perpendicular to the orbital plane, and $\dot{\theta}$ is the angular velocity of rotation. $A<B<C$ are the moments of inertia around the principal axis of the satellite around the $x$, $y$, $z$, respectively; a, $e$, $f(t)$, $r(t)=\frac{\textsf{a}(1-e^{2})}{1+e\cos f(t)}$, $n$ are the semi-major axis, orbital eccentricity, true anomaly, planet-satellite distance, and the mean-motion of the satellite, respectively.
It is well known that the dynamics of rotation of mid-sized satellites of the outer planets and the Moon are mainly characterized by synchronous orbit-rotation resonances (see Peale 1977, Melnikov and Shevchenko 2022). The exact synchronism occurs when the angular velocity of rotation of the satellite equals its mean motion, $n$. In practice, the synchronous resonance of the satellites is characterized by the physical libration of the angle $\Psi(t)=\theta(t)-f(t)$. $\Psi$ contains all main components of the spin-orbit perturbations, namely, the optical, free, and forced librations (see Callegari and Ribeiro 2015, and references therein). In the case of a small amplitude of libration of $\Psi$ and small orbital eccentricity of the satellite, the physics of the synchronous resonance can be interpreted by an analog of the forced harmonic oscillator (see Murray and Dermott 1999). Suppose also to model the shape of the satellite with a Roche ellipsoid with semi-axes $a$, $b$, and $c$ in the $x$, $y$, $z$ directions, respectively; in this case, $a>b>c$ (see Callegari and Rodríguez 2013, and references therein). Denote the free frequency by $\omega$ (which corresponds to the frequency of the oscillator in the harmonic approximation). The average linear theory [^1] W.r.t. mean anomaly and valid for small eccentricity. of synchronous resonance predicts that the ratio of the free and orbital frequencies $\frac{\omega}{n}$ is given by:
@@
The case of the Saturnian satellite Enceladus studied in Wisdom (2004) is a good example to illustrate the dynamics of secondary resonances. Under the hypothesis of a homogeneous body, and based on Voyager spacecraft data, the dimensions of Enceladus are $a=256.3\pm 0.3$ km, $b=247.3\pm 0.3$ km, $c=244.6\pm 0.3$ km (Dermott and Thomas 1994). From Equation (2) we obtain $\epsilon\sim 0.327$, a value close to $\frac{1}{3}$, so that Wisdom conjectured if Enceladus would be currently rotating close to the $\frac{\omega}{n}=\frac{1}{3}$ secondary resonance [^2] After Cassini images, the amplitude of the physical libration of Enceladus has been detected ($\sim 0.12$ degree; Thomas et al. 2016). This value is too high and therefore is not consistent with a homogeneous satellite. The best fitting given in Thomas et al. (2016) corresponds to a multi-layered satellite including a global subsurface ocean..
@@
Table 1: Selected physical and orbital parameters of the satellites studied in this work. In the case of Saturn, the values of the orbital eccentricities correspond to the maximum geometric values (Callegari et al. 2021;Callegari and Rodríguez 2023).
[…]
@@
The objects we study in this project belong to another class of even smaller inner satellites of the Jovian planets having mean diameters of a few kilometers or even less. We consider in this work the Saturnian satellites Methone (S/2004 S 1) and Aegaeon (S/2008 S 1) discovered by the Cassini spacecraft in 2004 (Porco 2004) and 2008 (Porco 2009), respectively. Similarly to Amalthea and Prometheus, their phase spaces also display the co-existence of the $\alpha$ and $\beta$-“resonances”. Methone, jointly with Aegaeon and Pallene is classified as ellipsoidal-like satellites [^3] The other denominations are irregular (e.g. Prometeus, Pandora, Epimetheus, Janus, Telesto, Calypso, and Helene), and irregular with equatorial ridges (e.g. Atlas, Pan, Daphnis) (Thomas and Helfenstein 2020).. We will focus next and throughout the paper on the case of Methone, and at the end of the paper, similar results will be shown and discussed for Aegaeon and Prometheus. The cases of Amalthea and Metis will also be revisited given the theory developed here.
@@
To investigate the rotation of Methone, we apply the standard model of spin-orbit resonance. We utilize the Everhart (1985) algorithm to solve numerically the rotation differential equations derived from hamiltonian (1). The full equation is non-autonomous such that, at first, we compute the surfaces of section (hereafter denoted by SOS), a practical procedure often applied in studies of rotation after Wisdom et al. (1984). In this technique, the pair of generalized variables $(\theta,\dot{\theta}/n)$ is plotted every time the satellite passes through the pericenter of its orbit. Figure 1a) shows the SOS for Methone in the vicinity of the synchronous domain. The fixed points of the $\alpha$ and $\beta$-“resonances” are indicated by vertical arrows. Note that the regimes are separated by a new separatrix (red curve in Fig. 1a)), while they are encompassed by a thin chaotic layer associated to the synchronous domain.
@@
The main goal of this work is to investigate the nature of the $\alpha$ and the $\beta$-regimes in the case of the satellite Methone. Both share the synchronous domain, so it is worth giving a cinematic description of the rotation of clones of Methone departing from initial conditions very close to their fixed points to identify their differences. Figs. 1(b,c) show the time variations of $\frac{\dot{\theta}}{n}$ and $\Psi=\theta-f$ corresponding to the $\alpha$ and $\beta$-regimes, respectively. The initial conditions are $\theta_{0}=0$ in both cases and $\frac{\dot{\theta}(0)}{n}=1$, $\frac{\dot{\theta}(0)}{n}=0.45$ in b) and c), respectively.
In the case of the $\beta$-regime (Fig. 1b)), $\Psi$ oscillates around zero with amplitude of $\sim 0.04$ radian (or $\sim 2$ degree) and $\frac{\dot{\theta}}{n}$ oscillates with amplitude of $\sim 0.036$. In both variables there is a “long-term” mode resulting from close values of the fundamental frequencies $n$ and $\omega$. In fact, from Equation (2) we have $\omega=n\epsilon$ and $\epsilon\sim 1.077$. Denoting by $T$, $P$, $T_{\omega}$ the periods of the beating, the orbital period, and the period of the free libration, we obtain $T\sim 13.2$ day from the relation $\frac{1}{T}=\frac{1}{P}-\frac{1}{T_{\omega}}$ considering $P\sim 1.014$ day and $T_{\omega}=P/\epsilon$.
@@
In the case of the $\alpha$-regime (Fig. 1c)), $\Psi$ oscillates around zero with amplitude of $\sim 0.5$ radian (or $\sim 29$ degree). $\frac{\dot{\theta}}{n}$ oscillates around the unit with amplitude of $\sim 0.5$ such that at the minima (indicated by black points in Fig. 1(b)), the satellite is always located at the pericenter of its orbit (open circles in Fig. 1(b)). Thus, due to this relatively large amplitude of oscillation of $\frac{\dot{\theta}}{n}$, the SOS displays $\frac{\dot{\theta}}{n}\sim 0.45$, the same as that of the initial value.
@@
Figure 1: (a) Surfaces of section of solutions of the model (1) applied to the satellite Methone. Each orbit has been integrated for 500 orbital periods of the satellite. The sections (projections) are made at each orbital revolution, and all the time the true anomaly of the satellite reaches $f=0$. The $x$ and $y$-axes are the (projected) rotation angle ($\theta$) around the axis of the largest moment of inertia and the ratio of the angular velocity of rotation and the mean motion, $\frac{\dot{\theta}}{n}$. The vertical arrows indicate the fixed points associated with the $\alpha$ and $\beta$ regimes of rotation. The numerical values of the parameters of the model adopted in this work which is necessary in the numerical solution are: $\textsf{a}=194,707.6$ km, $e=0.0014$, $a=1.94$ km, $b=1.29$ km (see Section 1 for definitions). (b,c) Time variations of $\frac{\dot{\theta}(t)}{n}$, $\Psi(t)=\theta(t)-f(t)$ for two trajectories within the domains of the $\alpha$ and $\beta$-regimes, respectively. $\theta_{0}=0$ in both cases and $\frac{\dot{\theta}(0)}{n}=1$, $\frac{\dot{\theta}(0)}{n}=0.45$ in b) and c), respectively. The horizontal dashed lines plotted in (c) guide the solutions close to the fixed points at the pericenter passage, while the horizontal line in (a) gives the straight $\frac{\dot{\theta}(t)}{n}=1$.
@@
Since the Methone’s phase space of rotation displays additional structures within the synchronous regime, all discussion given above leads us to investigate the existence of secondary resonances. Accordin to the most recent analyses of data taken from Cassini measurements, Thomas and Helfenstein (2020) determined that the dimensions of Methone are $a=1.94\pm 0.02$ km, $b=1.29\pm 0.04$ km, $c=1.21\pm 0.02$ km so that $\epsilon\sim 1.077$. This value of $\epsilon$ is close to the unit so that we conjecture if a 1:1 secondary resonance of the type $\frac{\omega}{n}\sim\frac{1}{1}$ would explain the existence of $\alpha$ and $\beta$ structures within the synchronous domain. To prove the conjecture we return to the Wisdom’s (2004) paper. He developed a simple analytical model for spin-orbit problem in hamiltonian formalism aiming to explain the phase space of Enceladus in the close vicinity of the synchronous regime. Having in hand an expanded version of hamiltonian of Equation (1), he selected the correct terms and explained analytically the properties of the secondary $\frac{1}{3}$ resonance. We follow his steps and reproduce its main results, and we choose those terms in hamiltonian which are proportional to the arguments containing the $\frac{1}{1}$ secondary resonance. By adopting a similar methodology given in Wisdom (2004), we successfully explain the co-existence of the $\alpha$ and $\beta$ in synchronous resonance by comparing the SOS with level curves of the constructed hamiltonian [^4] It is worth noting that the mapping of the 1:1 secondary resonance within the synchronous regime have been done in many works where very sophisticated models have been adopted (see Gkolias et al. 2016, Gkolias et al. 2019, Lei 2023.) .
@@
The presentation of the results of this paper is divided as follows. In Section 2, we deduce the expanded hamiltonian developed in Wisdom (2004) in detail, showing the terms of the hamiltonian explicitly (that ones related to equation 20 of the referred paper); some extensions of Wisdom’s model are also given. In Sections 3.1 and 3.2, we show exactly how the perturbed hamiltonian can give rise to the secondary resonances and write the hamiltonian for the $\frac{1}{1}$ secondary resonance. The level curves applied in the case of Methone are shown in Section 3.3. Analytical estimative of the bifurcations of the level curves are also provided. In Section 4 we analyze the cases of Aegaeon and Prometheus, and Section 5 is devoted to conclusions and more general discussion involving other satellites. Since the phase space of rotation of Methone, Aegaeon, Prometheus, Amalthea, Metis, and others share similar structures and properties within the synchronous regime, we are led to question their current states of rotation.
@@ 2.1 An integrable hamiltonian
The hamiltonian (1) is analogous to that one of a simple non-autonomous pendulum. An expanded version of the hamiltonian can be obtained resulting in a sum of autonomous-like pendulums. This can easily achieved after substituting in hamiltonian (1) the developed forms of $\left[\frac{\textsf{a}}{r(t)}\right]^{3}$, $\cos{2f(t)}$, $\sin{2f(t)}$:
@@
$\cos{2f(t)}\approx\cos{nt}+e(\cos{2nt}-1)+\frac{9}{8}e^{2}(\cos{3nt}-\cos{nt}),$ (3)
$\sin{2f(t)}\approx\sin{nt}+e\sin{2nt}+\frac{1}{8}e^{2}(9\sin{3nt}-7\sin{nt}).$
@@
where the term $-n\Phi$ comes from phase space extension (see Sussman and Wisdom 2001, Ferraz-Mello 2007). We have also neglected the term in $e^{2}$ since the orbital eccentricity of Methone is of the order of $10^{-3}$. Moreover, an irrelevant additive constant ($n^{2}C/2$) is not included in (7).
@@
$H^{{}^{\prime\prime}}_{0}(-,\tilde{J})$ $=$ $\frac{\epsilon^{2}n^{2}C}{4}\left(-1+\tilde{J}-\frac{1}{16}\tilde{J}^{2}-\frac{1}{256}\tilde{J}^{3}+\ldots\right),$ (8)
where
$\tilde{J}=\frac{4J}{\epsilon nC},$ (9)
such that
@@
$F_{1}$ $=$ $\left(2\tilde{J}+\frac{1}{4}\tilde{J}^{2}+\frac{27}{512}\tilde{J}^{3}+\ldots\right)^{1/2}\approx\sqrt{\tilde{J}}+\frac{1}{16}(\sqrt{2\tilde{J}})^{3},$ (12)
$F_{3}$ $=$ $\frac{1}{192}\left[(2\tilde{J})^{3}\left(1+\frac{9}{16}\tilde{J}+\ldots\right)\right]^{1/2}\approx\frac{1}{192}(\sqrt{2\tilde{J}})^{3}+\ldots.$ (13)
Equations (8)-(13) have been taken from Wisdom (2004). A proof of them can be obtained after applying the pendulum model to the hamiltonian (7) in the case of small oscillation approximation (see for instance appendix B.3 in Ferraz-Mello 2007).

### v3

# A Hamiltonian for 1/1 Rotational Secondary Resonances, and Application to Small Satellites of Saturn and Jupiter

@@ Abstract
In this work, we study the dynamics of rotation of the small satellites Methone and Aegaeon and revisit previous works on the rotation of Prometheus, Metis, and Amalthea. In all cases, the surfaces of section computed with the standard spin-orbit model reveal that the synchronous regime with small amplitude of libration shares another large domain in the phase space. We reproduce and apply the hamiltonian theory given in Wisdom (2004) to analytically characterize the detected structure as being a secondary resonance where the period of oscillation of the physical libration is similar to the orbital period of the satellite. We also show that the amplitude of libration around the secondary resonance is always larger than in the case of the other mode. Since the current rotational states of these sorts of satellites should be synchronous, our results can be considered in evolutionary studies of their rotation.
@@ 1 Introduction
where $\theta$ is an angle of rotation of the satellite measured from an inertial line; $p=C\dot{\theta}$ is the angular momentum of rotation around the $z-$axis assumed to be perpendicular to the orbital plane, and $\dot{\theta}$ is the angular velocity of rotation; $\epsilon^{2}=3\frac{B-A}{C}$, where $A<B<C$ are the moments of inertia around the principal axes of the satellite $x$, $y$, $z$, respectively; a, $e$, $f(t)$, $r(t)=\frac{\textsf{a}(1-e^{2})}{1+e\cos f(t)}$, $n$ are the semi-major axis, orbital eccentricity, true anomaly, planet-satellite distance, and the mean motion of the satellite, respectively.
It is well-known that the mid-sized satellites of the outer planets and the Moon rotate synchronously such that, roughly speaking, the velocity of angular rotation is equal to the mean motion (see Goldreich and Peale 1966, Peale 1977, Melnikov and Shevchenko 2022). The synchronous orbit-rotation resonance is characterized by the oscillation around zero of the angle $\Psi(t)=\theta(t)-f(t)$ present in Equation (1). $\Psi$ contains all main components of the spin-orbit perturbations, namely, the optical, free, and forced librations (see Murray and Dermott 1999, Callegari and Ribeiro 2015, and references therein). In the case of a small amplitude of libration of $\Psi$ and small orbital eccentricity of the satellite, the physics of the synchronous resonance can be interpreted by an analog of the forced harmonic oscillator (see Murray and Dermott 1999). Suppose also to model the shape of the satellite with a Roche ellipsoid with semi-axes $a$, $b$, and $c$ in the $x$, $y$, $z$ directions, respectively; in this case, $a>b>c$ (see Callegari and Rodríguez 2013, and references therein). Denote the free frequency by $\omega$ (corresponding to the frequency in the harmonic approximation). The average linear theory [^1] W.r.t. mean anomaly and valid for small eccentricity. of synchronous resonance predicts that the ratio of the free and orbital frequencies $\omega/n$ is given by:
@@
The case of the Saturnian satellite Enceladus studied in Wisdom (2004) is a good example to illustrate the dynamics of secondary resonances. Under the hypothesis of a homogeneous body, and based on Voyager spacecraft data, the dimensions of Enceladus were $a=256.3\pm 0.3$ km, $b=247.3\pm 0.3$ km, $c=244.6\pm 0.3$ km (Dermott and Thomas 1994). From Equation (2) we obtain $\epsilon\sim 0.327$, a value close to $\frac{1}{3}$, so that Wisdom conjectured if Enceladus would be currently rotating close to the $\frac{\omega}{n}=\frac{1}{3}$ secondary resonance [^2] After Cassini images, the amplitude of the physical libration of Enceladus has been detected ($\sim 0.12$ degree; Thomas et al. 2016). This value is too high and therefore is not consistent with a homogeneous satellite. The best fitting given in Thomas et al. (2016) corresponds to a multi-layered satellite including a global subsurface ocean. They also determine $a=256.2\pm 0.3$ km, $b=248.6\pm 0.2$ km, $c=252.24\pm 0.2$ km leading to $\epsilon\sim 0.300$..
@@
Table 1: Selected physical and orbital parameters of the satellites studied in this work. In the case of Saturn,
the values of the orbital eccentricities correspond to the maximum geometric values (Callegari et al. 2021;
Ceccatto et al. 2021, Callegari and Rodríguez 2023).
Methonea Aegaeona Prometheusa Amaltheab Metisb
@@
The objects we study in this project belong to another class of even smaller inner satellites of the Jovian planets having mean diameters of a few kilometers or even less. We consider in this work the Saturnian satellites Methone (S/2004 S 1) and Aegaeon (S/2008 S 1) discovered by the Cassini spacecraft in 2004 (Porco 2004) and 2008 (Porco 2009), respectively. Similarly to Amalthea and Prometheus, their phase spaces also display the co-existence of the $\alpha$ and $\beta$-“resonances”, as we will show. Methone, Aegaeon (and Pallene) are ellipsoidal-like satellites [^3] The other denominations are irregular (e.g. Prometeus, Pandora, Epimetheus, Janus, Telesto, Calypso, and Helene), and irregular with equatorial ridges (e.g. Atlas, Pan, Daphnis) (Thomas and Helfenstein 2020).. We will focus next and throughout the paper on the case of Methone, and at the end of the paper, similar results will be shown and discussed for Aegaeon and Prometheus. The cases of Amalthea and Metis will also be revisited given the theory developed here.
@@
To investigate the rotation of Methone, we apply the standard model of spin-orbit resonance. We utilize the Everhart (1985) algorithm to solve numerically the differential equations derived from hamiltonian (1). The full equation is non-autonomous such that, at first, we compute the surfaces of section (hereafter denoted by SOS), a practical procedure often applied in studies of rotation after Wisdom et al. (1984). In this technique, the pair of generalized variables $(\theta,\dot{\theta}/n)$ is plotted every time the satellite passes through the pericenter of its orbit. Fig. 1a) shows the SOS for Methone in the vicinity of the synchronous domain. The fixed points of the $\alpha$ and $\beta$-“resonances” are indicated by vertical arrows. Note that the regimes are separated by a new separatrix (red curve in Fig. 1a)), while they are encompassed by a thin chaotic layer associated to the synchronous domain.
@@
$\alpha$ and $\beta$-regimes share the synchronous domain, so it is worth giving a cinematic description of the rotation of clones of Methone departing from initial conditions very close to their fixed points to identify their differences. Figs. 1(b,c) show the time variations of $\frac{\dot{\theta}}{n}$ and $\Psi=\theta-f$ corresponding to the $\alpha$ and $\beta$-regimes, respectively. The initial conditions are $\theta_{0}=0$ in both cases, and $\frac{\dot{\theta}(0)}{n}=1.0211$, $\frac{\dot{\theta}(0)}{n}=0.45$ in b) and c), respectively.
@@
In the case of the $\beta$-regime (Fig. 1b)), $\Psi$ oscillates around zero with amplitude of $\sim 1$ degree and $\frac{\dot{\theta}}{n}$ oscillates with amplitude of $\sim 0.02$. In the case of the $\alpha$-regime (Fig. 1c)), $\Psi$ oscillates around zero with amplitude of $\sim 30$ degree. $\frac{\dot{\theta}}{n}$ oscillates around the unit with amplitude of $\sim 0.5$ such that at the minima, the satellite is always located at the pericenter of its orbit (full circles in Fig. 1c), bottom). Thus, due to this relatively large amplitude of oscillation of $\frac{\dot{\theta}}{n}$, the SOS displays $\frac{\dot{\theta}}{n}\sim 0.45$, the same as that of the initial value.
@@
Figure 1: (a) Surfaces of section of solutions of the model (1) applied to the satellite Methone. Each orbit has been integrated for 500 orbital periods of the satellite. The sections (projections) are made at each orbital revolution, and all the time the true anomaly of the satellite reaches $f=0$. The $x$ and $y$-axes are the (projected) rotation angle ($\theta$) around the axis of the largest moment of inertia and the ratio of the angular velocity of rotation and the mean motion, $\frac{\dot{\theta}}{n}$. The vertical arrows indicate the fixed points associated with the $\beta$ and $\alpha$-regimes of rotation. The horizontal line gives the straight $\frac{\dot{\theta}(t)}{n}=1$. The numerical values of the parameters of the model adopted in this work which are necessary in the numerical solution are: $\textsf{a}=194,707.6$ km, $e=0.0014$, $a=1.94$ km, $b=1.29$ km (see Section 1 for definitions). (b,c) Time variations of $\dot{\theta}(t)/n$, $\Psi(t)=\theta(t)-f(t)$ for two trajectories within the domains of the $\beta$ and $\alpha$-regimes, respectively. $\theta_{0}=0$ in both cases and $\frac{\dot{\theta}(0)}{n}=1.0211$, $\frac{\dot{\theta}(0)}{n}=0.45$ in b) and c), respectively. The horizontal dashed lines plotted in (b,c) guide the solutions close to the fixed points at the pericenter passage (indicated by full circles).
@@
Since the Methone’s phase space of rotation displays additional structures within the synchronous regime, all discussion given above leads us to investigate the existence of secondary resonances. According to the most recent analyses of data taken from Cassini measurements, Thomas and Helfenstein (2020) determined that the dimensions of Methone are $a=1.94\pm 0.02$ km, $b=1.29\pm 0.04$ km, $c=1.21\pm 0.02$ km so that $\epsilon\sim 1.077$. This value of $\epsilon$ is close to the unit so that we conjecture if a 1/1 secondary resonance of the type $\frac{\omega}{n}\sim\frac{1}{1}$ would explain the existence of $\alpha$ and $\beta$-structures within the synchronous domain. To prove the conjecture we return to the Wisdom’s (2004) paper. He developed a simple analytical model for spin-orbit problem in hamiltonian formalism aiming to explain the phase space of Enceladus in the close vicinity of the synchronous regime. Having in hand an expanded version of hamiltonian of Equation (1), he selected the correct terms and explained analytically the properties of the secondary 1/3 resonance. We follow his steps and reproduce its main results, and we choose those terms in hamiltonian which are proportional to the arguments containing the 1/1 secondary resonance. By adopting a similar methodology given in Wisdom (2004), we successfully explain the co-existence of the $\alpha$ and $\beta$-regimes in synchronous resonance by comparing the SOS with level curves of the constructed hamiltonian [^4] It is worth noting that the mapping of the 1/1 secondary resonance within the synchronous regime has been done in many works where very sophisticated models have been adopted (see Gkolias et al. 2016, Gkolias et al. 2019, Lei 2023.) .
@@
The presentation of the results of this paper is divided as follows. In Section 2, we deduce the expanded hamiltonian developed in Wisdom (2004) in detail, showing the terms of the hamiltonian explicitly (that ones related to equation 20 of the referred paper); some extensions of Wisdom’s model are also given. In Sections 3.1 and 3.2, we show exactly how the perturbed hamiltonian can give rise to the secondary resonances and write the hamiltonian for the 1/1 secondary resonance. The level curves applied in the case of Methone are shown in Section 3.3. Analytical estimation of the bifurcations of the level curves are provided in Section 3.4. In Section 3.5, we provide an equation to obtain the amplitude of oscillation at the fixed points of the $\alpha$ and $\beta$-regimes as a function of $\epsilon$; we show that the amplitude of oscillation around the secondary resonance is always larger than the case of the $\beta$-resonance. In Section 4, we analyze the cases of Aegaeon and Prometheus, and Section 5 is devoted to conclusions and more general discussion involving other satellites. Since the phase space of rotation of Methone, Aegaeon, Prometheus, Amalthea, Metis, and others share similar structures and properties within the synchronous regime, we are led to question their current states of rotation.
@@ 2.1 An integrable hamiltonian
The hamiltonian (1) is analogous to that one of a simple non-autonomous pendulum. An expanded version of the hamiltonian can be obtained resulting in a sum of autonomous-like pendulums. This can be easily achieved after substituting in hamiltonian (1) the developed forms of $\left[\frac{\textsf{a}}{r(t)}\right]^{3}$, $\cos{f(t)}$, $\sin{f(t)}$:
@@
$\cos{f(t)}\approx\cos{nt}+e(\cos{2nt}-1)+\frac{9}{8}e^{2}(\cos{3nt}-\cos{nt}),$ (3)
$\sin{f(t)}\approx\sin{nt}+e\sin{2nt}+\frac{1}{8}e^{2}(9\sin{3nt}-7\sin{nt}).$
@@
where the term $-n\Phi$ comes from phase space extension (see Sussman and Wisdom 2001, Ferraz-Mello 2007). We have also neglected the term in $25e^{2}/$ 2 since the orbital eccentricity of Methone is of the order of $10^{-3}$. Moreover, an irrelevant additive constant ($n^{2}C/2$) is not included in (7).
@@
$F_{1}$$\approx$ $\frac{2\sqrt{2}}{\lambda^{1/2}}J^{1/2}+\frac{\sqrt{2}}{2\lambda^{3/2}}J^{3/2}+\frac{23\sqrt{2}}{64\lambda^{5/2}}J^{5/2},$ (10)
$F_{3}$$\approx$ $\frac{\sqrt{2}}{12\lambda^{3/2}}J^{3/2}+\frac{3\sqrt{2}}{32\lambda^{5/2}}J^{5/2},$ (11)
where $\lambda=\epsilon nC$.
Equations (8)-(11) can be obtained from Wisdom (2004). Their deduction results of applying the pendulum model to the hamiltonian (7) in the case of small oscillation approximation (see for instance appendix B.3 in Ferraz-Mello 2007).